sadie5:

If a = 7, what is the value of the expression 2(a + 8)? A. 2 B. 15 C. 17 D. 23 E. 30

sadie5:

what do you think

supie:

Plug in the value of \(a\) into \(2(a + 8)\), \(a=7\) then use the distributive property \(a(b+c)=ab+ac\)

sadie5:

i think it's 30

supie:

Yes, correct

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sadie5 what do you think \(\color{#0cbb34}{\text{End of Quote}}\) how are you gonna ask urself what u think the answer is.-

Rigo:

lol lay